OpenStudy (anonymous):

3x +y = 12 ordered pair -2, ?

jimthompson5910 (jim_thompson5910):

Replace 'x' with -2

jimthompson5910 (jim_thompson5910):

3x+y = 12 3(-2)+y = 12 -6 + y = 12 Now solve for y

OpenStudy (anonymous):

y = 18?

jimthompson5910 (jim_thompson5910):

that's correct